### v2

@@ 3.2.6 Data Analysis Challenges and Solutions
Table 5: Summary of Data Analysis Challenges and Solutions
@@
Analyzing Large and Complex Data Solution referenced in Section 3.2.3 Proposed chen2016agile
Solution referenced in Section 3.2.4 Proposed dharmapal2016big
Solution referenced in Section 3.2.4 Proposed das2015towards
Solution referenced in Section 3.2.4 Proposed lin2018towards
@@
Analyzing Semantic Heterogeneity Data Solution referenced in Section 3.2.3 Proposed dos2021ontology
Solution referenced in Section 3.2.3 Implemented barcellos2020towards
Solution referenced in Section 3.2.3 Implemented rix2016agile
Efficient Data Analysis and Visualization Solution referenced in Section 3.2.3 Proposed dursun2014workflow
Solution referenced in Section 3.2.4 Proposed svensson2019unfulfilled
Solution referenced in Section 3.2.4 Proposed lehtonen2017visualizations
Solution referenced in Section 3.2.4 Proposed fabijan2016lack
Real-Time Data Analytics and Decision Making Solution referenced in Section 3.2.3, Section 3.2.4 Proposed chhillar2019act
Solution referenced in Section 3.2.4 Proposed olsson2018challenges
Selection of Appropriate Analytical Techniques Solution referenced in Section 3.2.3 Implemented harper2014agile
Solution referenced in Section 3.2.4 Proposed matthies2019towards
Solution referenced in Section 3.2.4 Proposed matthies2020playing
Solution referenced in Section 3.2.4 Proposed fagarasan2023integrating
Solution referenced in Section 3.2.4 Proposed batarseh2018predicting
Analyzing Large and Complex Data: Chen et al. chen2016agile delved into the challenge of developing a big data system that effectively supports advanced analytics by addressing the 5Vs of big data (Volume, Velocity, Variety, Veracity, and Value) and facilitating effective collaboration between data scientists and software engineers to maximize the value from big data. Dharmapal and Sikamani dharmapal2016big underscored the challenge of managing and processing the increased volume of unstructured data, which requires more complex enhancements in data management and analysis techniques. Das et al. das2015towards discussed the challenge of analyzing extensive volumes of data, particularly when the data is non-traditional or when the dataset is massive. Traditional analysis techniques often fall short due to these challenges. Lin et al. lin2018towards discussed the challenge of the high dimensionality of the datasets, which included detailed interactions and decisions made by the participants (students in the study). This complexity made it difficult to identify which features (which could include metrics such as the number of collaborators per task, the frequency of task updates, mood stability, and other behavioral indicators) or a combination of features could accurately predict certain behaviors of interest. Hamer et al. hamer2023students highlighted that interpreting the data provided by data-driven tools such as Git, Jira, and SonarQube can pose a significant barrier to beginning the effective use of these tools due to the large volume and complexity of the data generated. Furthermore, analyzing this data necessitates training to utilize the information correctly.
@@
Analyzing Semantic Heterogeneity Data: Dos Santos J’unior et al. dos2021ontology highlighted the difficulties organizations encounter when accessing, integrating, analyzing, and viewing data managed by disparate applications. Barcellos barcellos2020towards discussed the need for continuous software measurement, which involves collecting and analyzing data to provide useful information for daily activities and decision-making in software development. Rix et al. rix2016agile addressed the challenge of analyzing real-time data streams using data mining methods.
To address the analyzing semantic heterogeneity data challenges: The solution are the same as those of the data integration (as they co-occur together). We discussed those solutions in detail for data integration in Section 3.2.3.
Efficient Data Analysis and Visualization: Dursun et al. dursun2014workflow discussed the challenges of efficiently analyzing and visualizing vast and diverse oil and gas data and developing intelligent data-driven analytics software to optimize decision-making and minimize. Svensson et al. svensson2019unfulfilled addressed data analysis as a crucial step in data-driven decision-making (DDDM). The challenge lies in the fact that the quality of decisions is directly proportional to the quality of processing techniques and tools. In other words, ineffective or incorrect data analysis can lead to suboptimal or erroneous decisions. Lehtonen et al. lehtonen2017visualizations explored the utilization of visualizations for data analysis. The challenge lies in the potential complexity of these visualizations, which may hinder understanding. Fabijan et al. fabijan2016lack highlighted that companies are not capitalizing fully on the data they amass. This is attributed to challenges associated with efficiently and meaningfully integrating and analyzing customer data.
@@
Real-Time Data Analytics and Decision Making: Chhillar and Sharma chhillar2019act discussed using data analytics tools for generating metric reports. These reports are vital for indicating the overall health of the build post-deployment. The challenge here is the requirement for real-time data analytics to support decision-making processes. Olsson olsson2018challenges discussed the difficulty in handling and analyzing the post-deployment data collected from embedded systems.
@@
Selection of Appropriate Analytical Techniques: Harper and Dagnino harper2014agile discussed the need to specify analysis strategy, process substantial amounts of data, and complete analyses promptly. The challenge lies in fulfilling these expectations while managing the complexity inherent in advanced analytics applications. Matthies and Hesse matthies2019towards discussed data generation during software development processes such as work item descriptions, documentation, and version control information. The challenge lies in efficiently managing and analyzing this data. Matthies matthies2020playing discussed using various metrics designed for agile practices or tools that developers are already familiar with, such as the git command line, for project data analysis. The challenge is that improvement actions are often based on anecdotal evidence and experiences rather than empirical project evidence. Fagarasan et al. fagarasan2023integrating discussed the necessity for effective data analysis to extract meaningful insights from collected data. The challenge lies in interpreting, analyzing, and converting this information into actionable insights. Batarseh and Gonzalez batarseh2018predicting underscored the need to define variables for model development and investigate different models. The challenge here lies in selecting appropriate variables and models for analysis.
To address the selection of appropriate analytical techniques challenges: The solutions are the same as those of the data integration and collection (as they co-occur together). We discussed those solutions in detail in Sections 3.2.3 (data integration) and 3.2.4 (data collection).
@@ 3.3 Summary of Systematic Literature Review
Our SLR identified key data management challenges and solutions in agile software development. The review identified key data management aspects (with insights such as the overlapping between these aspects), including data integration, data quality, data collection, and data analysis. The SLR also revealed several solutions discussed to overcome those challenges. Figure 4 provides a summary of the challenges and links to solutions.
@@ 4 Practitioner Survey on Data Management
Now that we have identified data management challenges and solutions from the SLR in Section 3, we used those findings to design our practitioner survey. Based on the challenges and solutions reported in the literature, the survey aims to better understand what practitioners identify as challenges (RQ1) and solutions (RQ2) and whether those challenges and solutions are shared among practitioners. Furthermore, feedback from experts across different industries can enhance the relevance of the insights from the SLR for the practical application of our results and for guiding future research directions.
@@ 4.1.1 Survey Design
Target population and the sampling strategy: Our target population includes software development practitioners across various roles and experience levels in agile development projects. This encompasses senior developers, team leads, scrum masters, product owners, analysts, project managers, quality managers, DevOps engineers, data engineers, and solution architects. We employed a purposive sampling strategy patton2014qualitative , aimed at selecting individuals who are experienced in agile software development and with practical insights into data management challenges and solutions. This approach ensures that the data collected is rich in detail and relevant to our research questions. The online survey link was circulated to practitioners through personal connections and local networks. We aimed to distribute the survey to a broad audience, specifically those strongly interested in data management. The survey was distributed internationally to agile practitioners through direct contact at different organizations, through LinkedIn, and emails. The survey was distributed in March 2024, and the response period was open until June 2024. Responses were closely monitored to ensure the progress of the data collection process. To boost responses, follow-up reminders were sent to the contacts list to encourage participation.
@@
    Background: This section gathers demographics including age, gender, professional seniority, years of experience, educational background and role in agile projects, industry sector (pre-defined sectors with an option to add more) and professional information from participants. Additionally, we asked a mandatory question: “Have you worked, or are you currently working on a project that follows agile software development practices?”, If the answer was ’Yes,’ the participant could continue answering the subsequent questions.
@@
    Data management challenges and solutions: This section collects answers related to (data integration, collection, quality, and analysis) challenges and solutions based on the findings from the SLR (see Section 3), but also allowed participants to comment on their answers in an open-ended field. We employed a mix of multiple-choice and open-ended questions for each data management aspect section. For example, in the data integration section, we ask open-ended questions such as “What data integration challenges have you encountered in your projects?”, followed by multiple-choice to ask the same question with giving choices delivered from the SLR (see Appendix Table A1). The same style of questions (open-ended followed by multiple-choice) has been employed for each data management aspect section.
@@
Before deploying our survey, we conducted a pilot survey with three practitioners with more than ten years of experience in agile and data management (senior agile project manager, senior data analyst, and senior developer). The pilot aimed to verify the questions, scope, and estimated completion time. Feedback from the pilot was used to refine the questionnaire structure, clarify the questions, and address any identified validity threats. For example, we rephrased options for clarity and made structural changes to the survey flow. Additional questions were included based on the pilot results.
The survey can be accessed at https://forms.office.com/r/5b6vTcSjHV.
@@ 4.1.2 Data Analysis
We conducted a quantitative and qualitative analysis to analyze the survey data we obtained from all participants.
@@
  • Data Familiarising: we first familiarised ourselves with the data by reading over the survey responses. Iterative rereading of the responses helped to identify initial themes and codes. At each step, we took notes on possible codes and themes.
@@
  • Identifying Themes: codes were categorized into potential themes. Those themes were examined for consistency and coherence. Each theme was given a descriptive name that appropriately reflected its content We also identified sub-themes as needed. Two of the authors iteratively discussed and refined the identified themes. For instance, the codes “Inconsistent Data Structures” and “different Data Formats” were categorized under “Data Structure and Format Challenges”.
@@
      Level 1: Reviewing the Themes in Light of Participants’ Quotes Under the Theme: We checked the quotes from participants to make sure they accurately represented that theme.
@@
  • Comparison with SLR-Based Options: we compared the identified themes from the open-ended question with the options provided in the multiple-choice question. This helped integrate the findings into the results section.
@@ 4.2.1 Demographics
    Age: most participants are aged 35-44 (56.25%).
@@
    Industry Sectors: most participants primarily work in the ICT sector (44%). Other sectors (56%) include government administration, defence, public safety, education, manufacturing, public transport, healthcare, and retail.
@@ 4.2.2 Data Management Challenges and Solutions
Figure 6: Data Integration Challenges Reported by Practitioners
Data Integration Solutions: Participants have identified the use of communication-centric approaches as the most frequently adopted solution for data integration challenges (56%). Automated continuous testing is the least frequently mentioned solution (34%) (see Figure 7 and Table A5 in the Appendix). Participants also employed solutions such as customized technical solutions. For example, one participant noted that “custom scripts were developed to automate the transformation of data into the required formats and structures before loading them into the target system”. Another solution that was noted by practitioners is standardization and governance of data. Additionally, data segmentation and simplification (i.e., dividing complex data integration into smaller ones) was also mentioned as a solution.
Figure 7: Data Integration Solutions Reported by Practitioners
@@
Figure 8: Data Collection Challenges Reported by Practitioners
@@
Figure 9: Data Collection Solutions Reported by Practitioners
@@
Figure 10: Data Quality Challenges Reported by Practitioners
Data Quality Solutions: The participants have identified the use of automated data cleaning and enrichment tools (56%) and standardization of data quality metrics and processes (56%) as the most frequently adopted solutions for data quality challenges. Integration of advanced analytic platforms is the least reported solution (9%) (see Figure 11 and Table A7 in the Appendix). The participants also employed solutions such as training and awareness, which involves providing training on data quality best practices, and data restructuring, which involves modifying the structure of data to improve its quality.
Figure 11: Data Quality Solutions Reported by Practitioners
@@
Figure 12: Data Analysis Challenges Reported by Practitioners
Data Analysis Solutions: Compared to the findings from our SLR, the participants have identified the use of advanced analytical tools and techniques (50%) as the most frequently adopted solution for data analysis challenges. Adoption of machine learning and AI for data analysis (12.50%) is the least reported solution (see Figure 13 and Table A8 in the Appendix). The participants also highlighted that improve data integration, data quality and collection methods can facilitate the data analysis process.
Figure 13: Data Analysis Solutions Reported by Practitioners

### v4

@@ 3.2.7 Data Analysis Challenges and Solutions
Table 6: Summary of Data Analysis Challenges and Solutions identified by the SLR.
@@
Analyzing Large and Complex Data Ref in Sec. 3.2.4 Proposed chen2016agile
Ref in Sec. 3.2.5 Proposed dharmapal2016big
Ref in Sec. 3.2.5 Proposed das2015towards
Ref in Sec. 3.2.5 Proposed lin2018towards
@@
Analyzing Semantic Heterogeneity Data Ref in Sec. 3.2.4 Proposed dos2021ontology
Ref in Sec. 3.2.4 Implemented barcellos2020towards
Ref in Sec. 3.2.4 Implemented rix2016agile
Efficient Data Analysis and Visualization Ref in Sec. 3.2.4 Proposed dursun2014workflow
Ref in Sec. 3.2.5 Proposed svensson2019unfulfilled
Ref in Sec. 3.2.5 Proposed lehtonen2017visualizations
Ref in Sec. 3.2.5 Proposed fabijan2016lack
Real-Time Data Analytics and Decision Making Ref in Sec. 3.2.4 and Sec. 3.2.5 Proposed chhillar2019act
Ref in Sec. 3.2.5 Proposed olsson2018challenges
Selection of Appropriate Analytical Techniques Ref in Sec. 3.2.4 Implemented harper2014agile
Ref in Sec. 3.2.5 Proposed matthies2019towards
Ref in Sec. 3.2.5 Proposed matthies2020playing
Ref in Sec. 3.2.5 Proposed fagarasan2023integrating
Ref in Sec. 3.2.5 Proposed batarseh2018predicting
Analyzing Large and Complex Data: This challenge arises from the need to process and derive insights from massive, intricate datasets within agile environments. Agile teams often handle diverse data sources, requiring advanced tools and techniques to ensure timely and actionable analysis. Chen et al. chen2016agile delved into the challenge of developing an extensive data system that effectively supports advanced analytics by addressing the 5Vs of big data (Volume, Velocity, Variety, Veracity, and Value) and facilitating effective collaboration between data scientists and software engineers to maximize the value from big data. Dharmapal and Sikamani dharmapal2016big underscored the challenge of managing and processing the increased volume of unstructured data, which requires more complex enhancements in data management and analysis techniques. Das et al. das2015towards discussed the challenge of analyzing extensive volumes of data, particularly when the data is non-traditional or when the dataset is massive. Traditional analysis techniques often fall short due to these challenges. Lin et al. lin2018towards discussed the challenge of the high dimensionality of the datasets, which included detailed interactions and decisions made by the participants (students in the study). This complexity made it difficult to identify which features (which could include metrics such as the number of collaborators per task, the frequency of task updates, mood stability, and other behavioral indicators) or a combination of features could accurately predict certain behaviors of interest. Hamer et al. hamer2023students highlighted that interpreting the data provided by data-driven tools such as Git, Jira, and SonarQube can pose a significant barrier to beginning the effective use of these tools due to the large volume and complexity of the data generated. Furthermore, analyzing this data necessitates training to utilize the information correctly.
@@
Analyzing Semantic Heterogeneity Data: This challenge involves addressing inconsistencies in the meaning and interpretation of data across diverse systems and sources. Semantic heterogeneity complicates data integration and analysis in agile environments, where teams often rely on disparate tools and data formats. Dos Santos J’unior et al. dos2021ontology highlighted the difficulties organizations encounter when accessing, integrating, analyzing, and viewing data managed by disparate applications. Barcellos barcellos2020towards discussed the need for continuous software measurement, which involves collecting and analyzing data to provide useful information for daily activities and decision-making in software development. Rix et al. rix2016agile addressed the challenge of analyzing real-time data streams using data mining methods.
To address the analyzing semantic heterogeneity data challenges: The solutions are the same as those of the data integration (as they co-occur together). We discussed those solutions in detail for data integration in Section 3.2.4.
Efficient Data Analysis and Visualization: This refers to the challenge of transforming raw data into meaningful insights through streamlined analysis and clear visual representation. Agile projects demand rapid and effective visualization tools to aid in decision-making. Dursun et al. dursun2014workflow discussed the challenges of efficiently analyzing and visualizing vast and diverse oil and gas data and developing intelligent data-driven analytics software to optimize decision-making and minimize. Svensson et al. svensson2019unfulfilled addressed data analysis as a crucial step in data-driven decision-making (DDDM). The challenge lies in the fact that the quality of decisions is directly proportional to the quality of processing techniques and tools. In other words, ineffective or incorrect data analysis can lead to suboptimal or erroneous decisions. Lehtonen et al. lehtonen2017visualizations explored the utilization of visualizations for data analysis. The challenge lies in the potential complexity of these visualizations, which may hinder understanding. Fabijan et al. fabijan2016lack highlighted that companies are not capitalizing fully on the data they amass. This is attributed to challenges associated with efficiently and meaningfully integrating and analyzing customer data.
@@
Real-Time Data Analytics and Decision Making: This challenge focuses on enabling agile teams to process and utilize real-time data to support adaptive decision-making. Agile development’s iterative nature demands analytics systems that provide immediate feedback and insights. Chhillar and Sharma chhillar2019act discussed using data analytics tools for generating metric reports. These reports are vital for indicating the overall health of the build post-deployment. The challenge here is the requirement for real-time data analytics to support decision-making processes. Olsson olsson2018challenges discussed the difficulty in handling and analyzing the post-deployment data collected from embedded systems.
@@
Selection of Appropriate Analytical Techniques: This involves choosing the right tools, methods, and frameworks to analyze data effectively within agile contexts. With diverse data types and frequent changes, selecting appropriate techniques becomes critical to ensuring relevance and accuracy. Harper and Dagnino harper2014agile discussed the need to specify analysis strategy, process substantial amounts of data, and complete analyses promptly. The challenge lies in fulfilling these expectations while managing the complexity inherent in advanced analytics applications. Matthies and Hesse matthies2019towards discussed data generation during software development processes such as work item descriptions, documentation, and version control information. The challenge lies in efficiently managing and analyzing this data. Matthies matthies2020playing discussed using various metrics designed for agile practices or tools that developers are already familiar with, such as the git command line, for project data analysis. The challenge is that improvement actions are often based on anecdotal evidence and experiences rather than empirical project evidence. Fagarasan et al. fagarasan2023integrating discussed the necessity for effective data analysis to extract meaningful insights from collected data. The challenge lies in interpreting, analyzing, and converting this information into actionable insights. Batarseh and Gonzalez batarseh2018predicting underscored the need to define variables for model development and investigate different models. The challenge here lies in selecting appropriate variables and models for analysis.
To address the selection of appropriate analytical techniques challenges: The solutions are the same as those of the data integration and collection (as they co-occur together). We discussed those solutions in detail in Sections 3.2.4 (data integration) and 3.2.5 (data collection).
@@ 3.3 Summary of Systematic Literature Review
Our SLR identified key data management challenges and solutions in agile software development. The review identified key data management aspects (with insights such as the overlapping between these aspects), including data integration, data quality, data collection, and data analysis. The SLR also revealed several solutions discussed to overcome those challenges. Figure 4 summarizes the challenges and links to solutions.
@@ 4 Practitioner Survey on Data Management
Now that we have identified data management challenges and solutions from the SLR in Section 3, we used those findings to design our practitioner survey. Based on the challenges and solutions reported in the literature, the survey aims to better understand what practitioners identify as challenges (RQ1) and solutions (RQ2) and whether those challenges and solutions are shared among practitioners. Furthermore, feedback from practitioners across different industries can enhance the relevance of the insights from the SLR for the practical application of our results and for guiding future research directions.
@@ 4.1.1 Survey Design
Target population and the sampling strategy: Our target population includes software development practitioners across various roles and experience levels in agile development projects. This encompasses senior developers, team leads, scrum masters, product owners, analysts, project managers, quality managers, DevOps engineers, data engineers, and solution architects. We employed a purposive sampling strategy patton2014qualitative , aimed at selecting individuals experienced in agile software development and with practical insights into data management challenges and solutions. This approach ensures that the data collected is rich in detail and relevant to our research questions. We circulated the online survey link to practitioners through personal connections and local industrial networks. We aimed to distribute the survey to a broad audience, specifically those strongly interested in data management. The survey was distributed internationally to agile practitioners through direct contact at different organizations, through LinkedIn, and emails. The survey was distributed in March 2024, and the response period was open until June 2024. We closely monitored responses to ensure the progress of the data collection process. To boost the response rate, we sent follow-up reminders to the contacts list to encourage participation.
@@
    Background: This section gathers demographics including age, gender, professional seniority, years of experience, educational background and role in agile projects, industry sector (pre-defined sectors with an option to add more), and professional information from participants. Additionally, we asked a mandatory question: “Have you worked, or are you currently working on a project that follows agile software development practices?”; if the answer was ’Yes,’ the participant could continue answering the subsequent questions.
@@
    Data management challenges and solutions: This section collects answers related to (data integration, collection, quality, and analysis) challenges and solutions based on the findings from the SLR (see Section 3), but also allowed participants to comment on their answers in an open-ended field. We employed a mix of multiple-choice and open-ended questions for each data management aspect section. For example, in the data integration section, we ask open-ended questions such as “What data integration challenges have you encountered in your projects?”, followed by multiple-choice to ask the same question with giving choices delivered from the SLR (see Appendix Table A1). We employed the same style of questions (open-ended followed by multiple-choice) for each data management aspect section.
@@
Before deploying our survey, we conducted a pilot survey with three practitioners with more than ten years of experience in agile and data management (senior agile project manager, senior data analyst, and senior developer). The pilot aimed to verify the questions, scope, and estimated completion time. We used the feedback from the pilot to refine the questionnaire structure, clarify the questions, and address any identified validity threats. For example, we rephrased options for clarity and made structural changes to the survey flow. We included additional questions based on the pilot survey results.
The survey questionnaire can be accessed from our replication package fawzy_2024_10597818
@@ 4.1.2 Data Analysis
We conducted a quantitative and qualitative analysis of the survey data we obtained from all participants.
@@
  • Data Familiarizing: we first familiarized ourselves with the data by reading over the survey responses. Iterative rereading of the responses helped to identify initial themes and codes. At each step, we took notes on possible codes and themes.
@@
  • Identifying Themes: codes were categorized into potential themes. Those themes were examined for consistency and coherence. Each theme was given a descriptive name that appropriately reflected its content. We also identified sub-themes as needed. Two of the authors iteratively discussed and refined the identified themes. For instance, the codes “Inconsistent Data Structures” and “different Data Formats” were categorized under “Data Structure and Format Challenges”.
@@
      Level 1: Reviewing the themes in light of participants’ quotes under the theme: We checked participants’ quotes to ensure they accurately represented that theme.
@@
  • Comparison with SLR-Based Options: we compared the identified themes from the open-ended question with the options in the multiple-choice question. This helped integrate the findings into the results section.
@@ 4.2.1 Demographics
    Age: most participants are aged 35-44 (56%), 25–34 (25%), 45–54(16%), and 55–64 (3%).
@@
    Industry Sectors: most participants primarily work in the ICT sector (44%). Other sectors (56%) include government administration, defense, public safety, education, manufacturing, public transport, healthcare, and retail.
@@
Figure 5 shows an overview of the demographic of the survey participants. We can see that the survey participants come from a diverse composition of roles, industries, and experience levels, offering a broad perspective on data management challenges in agile development. We can see that tmostparticipants are senior-level professionals (62%), suggesting that the insights are informed by strategic and organizational viewpoints. Almost half (47%) of participants had over 10 years of professional expertise, representing mid-career professionals with significant industry experience. Most participants (94%) held tertiary degrees in computing-related fields and occupied various agile roles. We note that participants represent different industrial sectors, with 44% of participants from the ICT sector and the remainder from fields such as healthcare, education, public transport, and retail. This demographic composition highlights the study’s strength in capturing a broad spectrum of perspectives from practitioners across different industries. Although the survey is limited in its small sample size, the overrepresentation of participation from certain countries, and the predominance of specific roles, we made an effort to include a diverse range of industry sectors and seniority levels to overcome those limitations (discussed in more details in Section 6).
@@ 4.2.2 Data Management Challenges and Solutions
Figure 6: Distribution of Data Integration Challenges Reported by Practitioners
Data Integration Solutions: Participants have identified the use of
communication-centric approaches as the most frequently adopted solution for data integration challenges (56%). Automated continuous testing is the least frequently mentioned solution (34%) (see Figure 7 and Table A5 in the Appendix). Participants also employed solutions such as customized technical solutions. For example, one participant noted that “custom scripts were developed to automate the transformation of data into the required formats and structures before loading them into the target system”. Another solution that the practitioners noted is standardization and governance of data. Additionally, data segmentation and simplification (i.e., dividing complex data integration into smaller ones) was also mentioned as a solution.
Figure 7: Distribution of Data Integration Solutions Reported by Practitioners
@@
Figure 8: Distribution of Data Collection Challenges Reported by Practitioners
@@
Figure 9: Distribution of Data Collection Solutions Reported by Practitioners
@@
Figure 10: Distribution of Data Quality Challenges Reported by Practitioners
Figure 11: Distribution of Data Quality Solutions Reported by Practitioners
@@
Figure 12: Distribution of Data Analysis Challenges Reported by Practitioners
Data Analysis Solutions: Compared to the findings from our SLR, the participants have identified the use of advanced analytical tools and techniques (50%) as the most frequently adopted solution for data analysis challenges. Adoption of machine learning and AI for data analysis (12.50%) is the least reported solution (see Figure 13 and Table A8 in the Appendix). The participants also highlighted that improve data integration, data quality, and collection methods can facilitate the data analysis process.
Figure 13: Distribution of Data Analysis Solutions Reported by Practitioners
4.3 Alignment between the SLR and the Survey Findings
The survey findings confirm the challenges and solutions identified in the SLR across all four data aspects we examined (data integration, collection, quality, and analysis). For data integration, 97% (31 out of 32 participants) agreed with the SLR challenges. Additionally, 62.5% of participants identified managing the data integration process as a significant challenge. They (97%) supported solutions reported in the SLR, such as communication-centric methods (56%). Similarly, for data collection, 91% (29 out of 32 participants) agreed with the SLR data collection challenges, noting issues with capturing diverse data and automation (59%). Furthermore, 87.5% (29 participants) supported the proposed data collection solutions reported in the SLR, such as user-centered design strategies (59%). For data quality, 94% (30 participants) agreed with the data quality challenges reported in the SLR. They highlighted issues in ensuring accuracy and consistency (66%), while 84% (27 participants) agreed with the identified solutions, such as automated data cleaning and enrichment tools (56%). With regard to data analysis, 81% (26 participants) agreed with the data analysis challenges reported in the SLR. Specifically, 59% of participants reported difficulties with large datasets and real-time requirements. Additionally, 78% of the participants endorsed solutions such as advanced analytics tools (50%).
Overall, the majority of participants (91%) approved of the data management challenges we identified from the SLR. A large percentage of the participants (87%) also agreed with the proposed solutions noted in previous studies. These findings show a noteworthy alignment between the SLR and survey results, emphasizing the relevance of the identified challenges and the effectiveness of the proposed solutions.